PLEASE HELP WILL GIVE GREAT REVIEW!

Use what you need of the following set of information to answer the question:

S0 = $1.1420/€            S0 = £0.7776/$                        S0 = $1.2860/£       

                        E (S1 (3-months)) = $1.1280/€                              F3-months = $1.1590/€

                 Where: € = euro; £ = British pound; E ( ) = Expected.

  1. If you want to know the spot value of the euro in terms of the British pound from the information given above, how would you find out? Show calculations.

  1. If you want to know the spot value of the British pound in terms of the euro from the information given above, how would you find out? Show calculations
Answer #1

For calculating Euro in Terms of British Pound We need Dollar per Euro Value and Dollar per Pound Value.

Euro Per Dollar = 1/1.1420 = 0.8757

Dollar Per Pound = 1.2860

So Euro Per Pound = 0.8757 * 1.2860

= 1.1262 Approx

For Calculation Pound in terms of Euro we need Pound per Dollar Value and Dollar per Euro Value

Pound per Dollar = 0.7776

Dollar Per Euro = 1.1420

So Pound Per Euro = 0.7776 * 1.1420

= 0.8880 Approx